OpenStudy (1645323):

Which of the following polynomials has the lowest degree, a leading coefficient of 1, and -2 and 2 + 2i as roots? Select one: a. x3- 2x2+ 16 b. x3+ 2x2+ 16 c. x3- 2x2- 5x + 16 d. x3- 2x2- 16

OpenStudy (1645323):

its either c or d

OpenStudy (1645323):

@AnswerMyQuestions

OpenStudy (anonymous):

The correct answer is "a"

OpenStudy (anonymous):

x^3-2x^2+16=0 has a leading coefficient of 1 an the roots are: x1=-2 x2=2-2i x3=2+2i